OpenStudy (anonymous):

Other easy equation? 7a-(-b)=-6b+14

OpenStudy (anonymous):

@xxDragoonxx Help please?

OpenStudy (xxdragoonxx):

7a-(-b)=-6b+14 7a+b=-6b+14 7a=-6b+14-b = 7a=-7b+14 = a=-7b+14/7 = a =-7(b-2)/7 = a=-b+2

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

7a+7b=14

OpenStudy (anonymous):

@BoopStars is that the answer?

OpenStudy (anonymous):

Step 3: Divide both sides by 7.

OpenStudy (anonymous):

a1+b1=2?

OpenStudy (xxdragoonxx):

If solving it A = a=-b+2 If solving it B b=−a+2

OpenStudy (anonymous):

im so confused..

OpenStudy (anonymous):

1a-7b=14??? right???

Nnesha (nnesha):

do you have to solve for a , b or write in standard form

OpenStudy (anonymous):

solve for b

Nnesha (nnesha):

ok first of all distribute bracket by negative sign on left side

Nnesha (nnesha):

 now add 6b both side